### v1

# CITRAS: Covariate-Informed Transformer for Time Series Forecasting

Abstract
Covariates play an indispensable role in practical time series forecasting, offering rich context from the past and sometimes extending into the future. However, their availability varies depending on the scenario, and situations often involve multiple target variables simultaneously. Moreover, the cross-variate dependencies between them are multi-granular, with some covariates having a short-term impact on target variables and others showing long-term correlations. This heterogeneity and the intricate dependencies arising in covariate-informed forecasting present significant challenges to existing deep models. To address these issues, we propose CITRAS, a patch-based Transformer that flexibly leverages multiple targets and covariates covering both the past and the future forecasting horizon. While preserving the strong autoregressive capabilities of the canonical Transformer, CITRAS introduces two novel mechanisms in patch-wise cross-variate attention: Key-Value (KV) Shift and Attention Score Smoothing. KV Shift seamlessly incorporates future known covariates into the forecasting of target variables based on their concurrent dependencies. Additionally, Attention Score Smoothing transforms locally accurate patch-wise cross-variate dependencies into global variate-level dependencies by smoothing the past series of attention scores. Experimentally, CITRAS achieves state-of-the-art performance in both covariate-informed and multivariate forecasting, demonstrating its versatile ability to leverage cross-variate dependency for improved forecasting accuracy.
1 Introduction
Time series forecasting is a cornerstone in diverse fields such as energy [18, 35], retail [7, 5], and finance [6], where accurate predictions can drive strategic decisions and enhance operational efficiency. In practical scenarios, forecasters have access not only to the target variables they aim to predict but also to covariates that offer rich contexts for predictions. For instance, the electricity demand data shown in Figure 1 (Left) demonstrates a strong correlation with regional weather information and calendar data, highlighting the critical importance of appropriately incorporating covariates into the forecasting process.
Driven by the pressing demands of real-world challenges, this paper addresses time series forecasting that includes two types of covariates. The first is the observed covariate, which has recorded values from the past up to the point of prediction. The second is the known covariate, which has values that are known from the past through to the future forecasting horizon, such as calendar events and promotion schedules.
Figure 1: Left: Hourly electricity demand data for three weeks in Panama [2]. Temperature and humidity generally show a positive correlation with demand but exhibit local disturbances, as indicated by the red circle. Holiday and weekend indicators demonstrate a negative correlation with demand, providing important contextual information in future forecasting horizon. Right: Performance improvement of CITRAS through the utilization of past (observed and past known) covariates and future (future known) covariates.
This practical problem setting presents two significant challenges. First, the forecasting model must be capable of universally accommodating heterogeneous variables. The availability and number of covariates can vary depending on the scenario, and situations often involve multiple target variables simultaneously, which is referred to as multivariate forecasting. Second, the model must capture the dependencies between variables from both fine-grained and coarse-grained perspectives. The fine-grained perspective enables the model to precisely capture the local impact of covariates, such as sudden drops in electricity demand on holidays. However, models that rely solely on a fine-grained perspective may overlook global dependencies between variables, as some covariates may be infrequent (e.g., holidays, promotions) or exhibit temporal disturbances. Therefore, it is crucial for the model to capture dependencies between variables at multiple levels to ensure robust and accurate forecasting.
In recent years, deep forecasting models [33], particularly Transformers [32], have made remarkable progress. However, many studies have primarily focused on the dependencies among target variables. Some pioneering works [34, 24] have highlighted the importance of covariates; however, the effective integration of covariates, especially future known covariates, requires further exploration. Existing approaches to capture cross-variate dependencies can be broadly categorized into two types: variate-level and patch-level. iTransformer [22] is a representative variate-level approach that captures inter-variate correlations by embedding each sequence into a single variate token and applying attention across multiple variate tokens. On the other hand, patch-level approaches, represented by Crossformer [43], aim to capture more fine-grained cross-variate dependencies by applying attention between locally semantic tokens obtained through patching. However, these methods are polarizing, and an architecture capable of capturing both global and local dependencies has yet to be developed.
To achieve effective time series forecasting that captures the intricate dependencies among diverse variables, we propose CITRAS: Covariate-Informed Transformer for time series forecasting. This model is a patch-based transformer that has a cross-time attention module and a cross-variate attention module separately. To effectively model the multi-granular cross-variate dependency, we introduce Attention Score Smoothing in the cross-variate attention module, which transforms locally accurate patch-wise dependencies into global variate-level dependencies. Furthermore, we introduce Key-Value (KV) Shift mechanism, which associates the key of the known covariate with the value from one patch step ahead, thereby seamlessly integrating future known covariates into the prediction process. These innovations expand the covariate understanding capabilities of the canonical Transformer, fully preserving its strong autoregressive properties.
In summary, our contributions are as follows:
  • To fully utilize the rich context offered by covariates, we extend the Transformer to flexibly accommodate observed covariates and known covariates.
  • Technically, we introduce Attention Score Smoothing and KV Shift into the cross-variate attention module. These enhancements unravel the multi-granular cross-variate dependencies inherent in time series data and enable the effective utilization of future known covariates in predictions.
  • Our resulting model, CITRAS, achieves state-of-the-art performance on both covariate-informed settings and multivariate settings, demonstrating its versatile ability to leverage cross-variate and cross-time dependency for improved forecasting accuracy.
2 Related Work
2.1 Transformer-based Time Series Forecasting
Given the significant success of Transformers in the fields of natural language processing [10] and computer vision [12], numerous studies have attempted to leverage their capabilities for time series forecasting. These works can be roughly divided into two approaches: channel-dependent (CD) and channel-independent (CI). The CD approach assumes that the future values of a variable are determined by its past values as well as the values of other variables. In contrast, the CI approach posits that a specific variable only depends on its past values, omitting interactions between variables.
Early applications of Transformers in time series forecasting can be viewed as CD approaches, where multiple variables at the same time point are embedded into temporal tokens [44, 40, 45, 20, 11]. To address the computational costs and intricate dependencies arising from long time series data, Informer [44] designs a ProbSparse self-attention mechanism with a distilling operation to efficiently focus on important features. Autoformer [40] incorporates decomposition and an auto-correlation mechanism to uncover reliable dependencies from complex temporal patterns. FedFormer [45] employs frequency-domain sparse representations to efficiently capture the global properties of time series. However, point-wise temporal tokens have limited local semantics, making it difficult to capture intricate cross-time and cross-variate dependencies.
In response to these challenges, PatchTST [27] adopts a CI approach and introduces patching, which handles time series data by segmenting it into fixed-length segments. Patching reduces computational costs and enhances the local semantics in token representation, making it a prevalent technique in subsequent methods [42, 41, 47]. Additionally, the CI approach is also adopted by many recent large models [25, 3, 9, 13] and LLM-based forecasters [14, 46, 15], as it avoids the risk of unnecessary noisy interactions between channels.
However, the CI approach is sometimes regarded as an oversimplification [1], especially considering that fluctuations in time series data are often influenced by external factors. A representative CD approach is iTransformer [22], which embeds each channel into a single variate token and applies cross-variate attention over them. TimeXer [34] also captures cross-variate dependencies at the variate level, with the additional use of cross-time attention to capture intra-variate temporal dependencies. Meanwhile, Crossformer [43] applies cross-variate attention to patch-level representations. Recent Any-variate Attention techniques [37, 24, 21] also apply attention to patch-level representations obtained by flattening all variables into one sequence. While the patch-level approach captures more fine-grained dependencies between variables, its local receptive field lacks global context awareness. Our method, CITRAS, first captures fine-grained dependencies at the patch level, and Attention Score Smoothing gradually refines them to achieve global variate-level dependencies.
2.2 Covariate-Informed Time Series Forecasting
To leverage the rich context offered by covariates, statistical models such as ARIMAX [36] and VARX [26] extend the original models by assuming linear relationships between variables. To capture more intricate dependencies through deep learning, N-BEATSx [28] enhances the original N-BEATS [29] by introducing an additional exogenous stack. TiDE [8] proposes an MLP-based architecture that utilizes future known covariates in the final projection to leverage their direct effects on future target variables. Temporal Fusion Transformer [19] designs gating mechanisms on top of the Transformer to select relevant features for forecasting. While these approaches adopt elaborate model designs to capture covariates, recent methods, such as TimeXer [34] and Timer-XL [24], aim to integrate covariates by leveraging the inherent expressive power of the Transformer. We follow this line of work and extend the Transformer to accommodate covariates while fully preserving the original architecture. Additionally, we focus primarily on the utilization of future known covariates, which, despite their importance, have not been sufficiently discussed in these studies.
3 CITRAS
3.1 Problem setting
Let $\mathbf{X}_{1:T}^{tgt,:}=\{\mathbf{X}_{1:T}^{tgt,1},\mathbf{X}_{1:T}^{tgt,2},...,\mathbf{X}_{1:T}^{tgt,C_{tgt}}\}\in\mathbb{R}^{T\times C_{tgt}}$ be a target time series of length $T$ and a number of variables $C_{tgt}$. With optional usage of observed covariates $\mathbf{X}_{1:T}^{obs,;}\in\mathbb{R}^{T\times C_{obs}}$ and known covariates $\mathbf{X}_{1:T+S}^{knw,:}\in\mathbb{R}^{(T+S)\times C_{knw}}$, the goal of the forecasting model $\mathcal{F}_{\theta}$ is to predict the future $S$ time steps of the target time series, which can be formulated as follows:
$\widehat{\mathbf{X}}_{T+1:T+S}^{tgt,:}=\mathcal{F}_{\theta}\left(\mathbf{X}_{1:T}^{tgt,:},\mathbf{X}_{1:T}^{obs,:},\mathbf{X}_{1:T+S}^{knw,:}\right)$ (1)
3.2 Architecture
As shown in Figure 2, CITRAS is a patch-based, decoder-only transformer with $L$ layers of cross-time attention modules and cross-variate attention modules. Our core mechanisms, KV Shift and Attention Score Smoothing, reside in the cross-variate attention module.
Figure 2: Overall structure of CITRAS. (a) The embedding module applies patching to each variable, yielding temporal token embeddings. (b) The Cross-Time Attention module captures intra-variate cross-time dependencies. (c) The Cross-Variate Attention module captures cross-variate dependencies. (d) KV Shift associates the key of the known covariate with the value from one patch step ahead. Attention Score Smoothing calculates the Exponential Moving Average (EMA) of patch-wise attention scores. (e) The projection layer projects each target token embedding to the values of the next patch.
Embedding
First, all sequences are segmented into non-overlapping patches of length $P$, and each patch is embedded into a $D$-dimensional token. Taking target variables as an example, this can be formalized as:
$\{\mathbf{s}_{1}^{tgt,c},\mathbf{s}_{2}^{tgt,c},...,\mathbf{s}_{N_{tgt}}^{tgt,c}\}$ $=\operatorname{Patchfy}\left(\mathbf{X}_{1:T}^{tgt,c}\right),~{}~{}$ $c=1,...,C_{tgt}$ (2)
$\mathbf{H}_{i}^{tgt,c}$ $=\operatorname{Embed}\left(\mathbf{s}_{i}^{tgt,c}\right),~{}~{}$ $i=1,...,N_{tgt},~{}~{}c=1,...,C_{tgt}$
where $N_{tgt}=\lfloor\frac{T}{P}\rfloor$ is the number of patches and $\operatorname{Embed}:\mathbb{R}^{P}\to\mathbb{R}^{D}$ is a shared linear projector across all variables. We denote the token embeddings of a target variable at all patch steps as $\mathbf{H}_{:}^{tgt,c}=~{}\{\mathbf{H}_{i}^{tgt,c}\}_{i=1}^{N_{tgt}}~{}\in~{}\mathbb{R}^{N_{tgt}\times D}$. Similarly, the token embeddings of an observed covariate are denoted as $\mathbf{H}_{:}^{obs,c}\in\mathbb{R}^{N_{obs}\times D}$, and those of a known covariate are denoted as $\mathbf{H}_{:}^{knw,c}\in\mathbb{R}^{N_{knw}\times D}$, all obtained using the shared parameters with the target variables. Here, $N_{obs}=N_{tgt}=\lfloor\frac{T}{P}\rfloor$ and $N_{knw}=\lfloor\frac{(T+S)}{P}\rfloor$.
Cross-Time Attention
In the cross-time attention module, we apply multi-head attention with causal masking to all variables to capture their intra-variate cross-time dependency. Here, we adopt Rotary Position Embedding (RoPE) [31] to capture temporal order. Taking target variables as an example, and dropping layer index for brevity, this can be formalized as:
$\widetilde{\mathbf{H}}_{:}^{tgt,c}$ $=\operatorname{LN}\left(\mathbf{H}_{:}^{tgt,c}+\operatorname{MHA}\left(\mathbf{H}_{:}^{tgt,c},\mathbf{H}_{:}^{tgt,c},\mathbf{H}_{:}^{tgt,c}\right)\right),~{}~{}$ $c=1,...,C_{tgt}$ (3)
$\mathbf{H}_{:}^{tgt,c}$ $=\operatorname{LN}\left(\widetilde{\mathbf{H}}_{:}^{tgt,c}+\operatorname{FFN}\left(\widetilde{\mathbf{H}}_{:}^{tgt,c}\right)\right),~{}~{}$ $c=1,...,C_{tgt}$
where $\operatorname{LN}$ denotes layer normalization [4], $\operatorname{MHA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)$ denotes the multi-head attention layer where $\mathbf{Q}$,$\mathbf{K}$, and $\mathbf{V}$ serve as queries, keys and values, and $\operatorname{FFN}$ denotes a feed-forward network.
Again, token embeddings of observed covariates $\mathbf{H}_{:}^{obs,c}$ and those of known covariates $\mathbf{H}_{:}^{knw,c}$ are similarly processed with the shared parameters.
Cross-Variate Attention — KV Shift
Attention is recognized as an effective approach for capturing cross-variate dependencies [22]. To represent the dependencies of target variables while avoiding unnecessary interactions among covariates, we adopt multi-head attention, where queries originate from the target variables, whereas keys and values are derived from all available variables. However, applying attention across the variables of the same patch steps, as common in [43, 1], fails to leverage pivotal future patches of known covariates, as they do not have corresponding target patches.
To address this issue, we introduce KV Shift mechanism in the multi-head attention layer of the cross-variate attention module, which associates the key of the known covariate with the value from one patch step ahead. Specifically, the key and value at patch step $i$ can be formalized as:
[…]
$\widetilde{\mathbf{H}}_{i}^{tgt,:}$ $=\operatorname{LN}\left(\mathbf{H}_{i}^{tgt,:}+\operatorname{MHA}\left(\mathbf{H}_{i}^{tgt,:},\mathbf{H}_{i}^{k,:},\mathbf{H}_{i}^{v,:}\right)\right),~{}~{}$ $i=1,...,N_{tgt}$ (5)
$\mathbf{H}_{i}^{tgt,:}$ $=\operatorname{LN}\left(\widetilde{\mathbf{H}}_{i}^{tgt,:}+\operatorname{FFN}\left(\widetilde{\mathbf{H}}_{i}^{tgt,:}\right)\right),~{}~{}$ $i=1,...,N_{tgt}$
Intuitively, KV Shift allows the target variable (query) to incorporate the next-step future known covariate (value) based on the concurrent dependency on that covariate (key). As the target token gradually transforms into the prediction for the next step, this facilitates a natural flow of information for leveraging future known covariates.
Cross-Variate Attention — Attention Score Smoothing
The attention score of $\mathrm{MHA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)$ can be calculated as:
$\widetilde{\mathbf{A}}=\left(\mathbf{Q}\mathbf{W}_{q}\right)\left(\mathbf{K}\mathbf{W}_{k}\right)^{\top}$ (6)
where $\mathbf{W}_{q},\mathbf{W}_{k}\in\mathbb{R}^{D\times d_{k}}$ and $d_{k}$ is the dimension of the query, key, and value. In the cross-variate attention module, this attention score is calculated for each patch step separately, capturing fine-grained dependency between variables at that step. However, time series data often exhibit local disturbances, and sparse variables may maintain constant values within a patch. In such scenarios, local patch-wise attention fails to represent global variate-level dependencies.
To address this issue, we introduce Attention Score Smoothing in the multi-head attention layer of the cross-variate attention module, which transforms locally accurate patch-wise dependencies into global variate-level dependencies. Specifically, the attention score at each step is smoothed based on the attention scores up to that step in each head. We employ an Exponential Moving Average (EMA) as a smoothing method, as it enables adaptation to shifting correlations between variables [16] by assigning exponentially decreasing weights over time. Denoting original attention score at patch step $i$ as $\widetilde{\mathbf{A}}_{i}$, the smoothed attention score $\mathbf{A}_{i}$ can be calculated as:
$\mathbf{A}_{i}$ $=\alpha\widetilde{\mathbf{A}}_{i}+\left(1-\alpha\right)\mathbf{A}_{i-1},~{}~{}i=2,...,N_{tgt}$ (7)
where $\mathbf{A}_{1}=\widetilde{\mathbf{A}}_{1}$ and $\alpha$ is a smoothing factor, which we consider as a shared hyperparameter across all cross-variate attention modules and heads. Note that the aim and structure of Attention Score Smoothing are completely different from the exponential smoothing attention in ETSformer [38], which adopts an exponentially decreasing weights based on the time lag instead of the vanilla attention scores to capture intra-variate cross-time dependency.
Projection
Following the next token prediction approach common in decoder-only Transformers, each token embedding of the target variables is used to predict the values of the next patch. For $i=1,...,N_{tgt}$ and $c=1,...,C_{tgt}$, this can be formalized as:
$\widehat{\mathbf{X}}_{iP+1:(i+1)P}^{tgt,c}=\operatorname{Project}\left(\mathbf{H}_{i}^{tgt,c}\right)$ (8)
where $\operatorname{Project}:\mathbb{R}^{D}\to\mathbb{R}^{P}$ is a shared linear projector across all steps and target variables. In the training phase, we calculate the squared loss using all of these outputs.
4 Experiments
To verify the effectiveness and versatility of CITRAS, we extensively evaluated it in the two settings, short-term covariate-informed forecasting and well-acknowledged long-term multivariate forecasting.
4.1 Covariate-Informed Forecasting
Dataset
For covariate-informed forecasting, we use the short-term electricity price forecasting (EPF) dataset [18] and electricity demand forecasting (EDF) dataset [2]. The EPF dataset consists of six years of hourly time series data from five major power markets, with electricity price as the target and day-ahead load forecasts and generation forecasts as known covariates. The EDF dataset is an hourly time series dataset spanning over five years, where the target variable is the national electricity demand in Panama. We adopt four meteorological indices measured in three cities (total 12 indices) as observed covariates, and three binary calendar indicators (holiday, weekend, school day) as known covariates. Following previous works [28], we set the input length as 168, forecasting length as 24, and patch length as 24. Further details about the datasets are included in the Appendix A.1.
Baselines
We include eight strong deep forecasting models, namely Timer-XL[24], TimeXer [34], iTransformer [22], TiDE [8], Temporal Fusion Transformer (TFT) [19], Crossformer [43], TimesNet [39], and Autoformer [40]. Notably, TiDE and TFT are originally developed to accommodate future known covariates. For Timer-XL, TimeXer and iTransformer, we extend their original implementation to support future known covariate, as detailed in the Appendix A.3. Crossformer, TimesNet and Autoformer remain original implementation and cannot accommodate future known covariates.
Result
Comprehensive results for covariate-informed forecasting are presented in Table 1. In the EPF dataset, the day-ahead load forecasts and generation forecasts show a strong correlation with future electricity prices, providing valuable insights for forecasting. Crossformer, TimesNet, and Autoformer lack the ability to capture these known covariates, resulting in inferior performance. iTransformer and TimeXer capture these covariates at the variate level, while Timer-XL captures them at the patch level. These contrasting approaches show competitive performance. On the other hand, CITRAS incorporates these covariates at a locally accurate patch level and captures dependencies at the variate level as well, demonstrating superior accuracy. The EDF dataset poses further challenges to the models due to the presence of numerous observed and known covariates, some of which exhibit local disturbances. Interestingly, Timer-XL, which applies attention to all token embeddings at once, performs worse than other variate-level approaches due to noisy interactions. In contrast, CITRAS applies separate attention to cross-time and cross-variate directions and employs Attention Score Smoothing, allowing it to robustly capture dependencies despite disturbances, thereby outperforming other baselines.
Table 1: Full results of the covariate-informed forecasting, where future known covariates are available. We adhere to the standard protocol [28], setting the input length and forecasting length to 168 and 24, respectively. Avg represents the average results from all EPF datasets. EPF results for Crossformer, TimesNet, and Autoformer are cited from [24], while others are reproduced. An asterisk (*) indicates that the model cannot accommodate future known covariates.
Model
CITRAS
Timer-XL
TimeXer
iTrans.
TiDE
TFT
Cross.*
TimesNet*
Auto.*
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
EPF
NP
0.171
0.214
0.176
0.220 0.175
0.225
0.203
0.250
0.296
0.309
0.193
0.237
0.240
0.285
0.250
0.289
0.402
0.398
PJM
0.065
0.152
0.062
0.155
0.080
0.175
0.077
0.171
0.105
0.213
0.079
0.174
0.101
0.199
0.097
0.195
0.168
0.267
BE
0.369
0.251
0.384
0.257
0.368
0.231
0.339
0.246
0.447
0.301
0.352 0.240
0.420
0.290
0.419
0.288
0.500
0.333
FR
0.364
0.180
0.430
0.214
0.364
0.193
0.378
0.205
0.407
0.251
0.371
0.182
0.434
0.208
0.431
0.234
0.519
0.295
DE
0.217
0.293
0.228 0.299
0.283
0.322
0.258
0.327
0.515
0.462
0.262
0.326
0.574
0.430
0.502
0.446
0.674
0.544
Avg
0.237
0.218
0.256
0.229
0.254
0.229 0.251
0.240
0.354
0.307
0.251
0.232
0.354
0.284
0.340
0.290
0.453
0.368
EDF
0.070
0.193
0.079
0.206
0.071
0.199
0.071 0.197
0.133
0.259
0.092
0.225
0.108
0.233
0.101
0.233
0.193
0.327
4.2 Multivariate forecasting
For a comprehensive comparison, we further evaluate our model on well-recognized long-term multivariate forecasting datasets. To reflect practical scenarios, we adopt the rolling forecasting approach [23], where a model trained with predetermined input and forecasting lengths is evaluated with a longer forecasting length by iteratively integrating its output into the subsequent input. Following previous works [23, 24], we set the input length to 672, the forecasting length to {96, 192, 336, 720}, and the patch length to 96. We present the average results for the four forecasting lengths in Table 2. CITRAS ranks first or second in most benchmarks, highlighting the effectiveness and versatility of its design.
Table 2: Results of the multivariate forecasting, where the input length is set to 672, and the forecasting length is set to {96, 192, 336, 720}. We adopt the rolling forecasting approach following [23] where one model is used for four forecasting lengths. Averaged results are reported here and full results are provided in Table 5. Results are cited from Timer-XL [24] if available; otherwise reproduced.
Model
CITRAS
Timer-XL
TimeXer
iTrans.
TFT
Cross.
TimesNet
Auto.
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTh1
0.389
0.421
0.409 0.430
0.428
0.453
0.421
0.445
0.603
0.556
0.448
0.468
0.495
0.491
0.517
0.528
ETTh2
0.359
0.405
0.372
0.409
0.353
0.405
0.383
0.415
0.405
0.440
0.814
0.649
0.362
0.412
0.397
0.442
ETTm1
0.350
0.382
0.368 0.390 0.368
0.394
0.377
0.404
0.541
0.498
0.368
0.399
0.379
0.406
0.550
0.505
ETTm2
0.262
0.320
0.275
0.327
0.266 0.321
0.278
0.334
0.389
0.399
0.396
0.442
0.303
0.347
0.338
0.393
ECL
0.160 0.253
0.155
0.246
0.166
0.264
0.164
0.258
0.199
0.300
0.168
0.267
0.201
0.303
0.289
0.379
Weather
0.233
0.272
0.240
0.273
0.223
0.263
0.266
0.291
0.308
0.333
0.225
0.286
0.264
0.293
0.435
0.455
Traffic
0.379
0.254
0.374
0.255
0.401
0.288
0.384
0.274
0.590
0.322
0.513
0.294
0.602
0.322
0.684
0.433
4.3 Ablation Study
Table 3: Ablation Results of Attention Score Smoothing
Setting
w/o
w/
Metric
MSE
MSE
EPF-NP
0.175
0.171 (-2.8%)
EDF
0.074
0.070 (-4.3%)
ETTh1
0.401
0.389 (-3.2%)
ECL
0.165
0.160 (-3.2%)
CITRAS adopts two novel mechanisms: KV Shift and Attention Score Smoothing. To clarify the effectiveness of KV Shift, as well as the importance of covariates, we trained and tested our model under three covariate usage settings: None (no covariates used), P (observed and past known covariates used), and P+F (P and future known covariates used). Figure 3 demonstrates that forecasting solely relying on the target has limitations, and covariates play a crucial role in performance improvement. In particular, future known covariates provide much richer context extending into the forecasting horizon, highlighting the indispensability of KV Shift to incorporate them. Furthermore, compared to other baselines, CITRAS shows a higher rate of performance improvement through the use of covariates, supporting its suitability for covariate-informed forecasting.
We also trained and tested an ablated version of our model without Attention Score Smoothing. Table 3 shows that Attention Score Smoothing is effective in both covariate-informed and multivariate settings. This supports the versatility of this mechanism, which captures variate-level cross-variate dependencies by smoothing over detailed patch-level dependencies.
Figure 3: Comparison of MSE on the EPF-NP and EDF datasets under three covariate usage settings: None (no covariates used), P (observed and past known covariates used), and P+F (P and future known covariates used). The MSE change rates shown on the bar graph are calculated with respect to the None setting as the baseline.
5 Conclusion
Considering the critical roles of covariates in practical time series forecasting, we propose CITRAS, a patch-based Transformer that flexibly leverages multiple targets, observed covariates, and known covariates. Specifically, we introduce two novel mechanisms within the patch-wise cross-variate attention modules: Key-Value (KV) Shift, which seamlessly integrates future known covariates with other variables, and Attention Score Smoothing, which transforms locally accurate dependencies into global variate-level dependencies. Experimentally, CITRAS achieves state-of-the-art performance and demonstrates strong cross-variate understanding in both covariate-informed and multivariate settings. In the future, we will explore foundational temporal understanding capability in covariate-informed settings.
Appendix A Implementation Details
A.1 Dataset Descriptions
We perform short-term covariate-informed forecasting on the electricity price forecasting (EPF) datasets and the electricity demand forecasting (EDF) dataset. The EPF datasets [18] consist of five subsets from different day-ahead electricity markets, each spanning six years: (1) EPF-NP documents the Nord Pool electricity market from 2013-01-01 to 2018-12-24, containing hourly electricity prices as a target variable, with corresponding grid load and wind power forecasts as known covariates. (2) EPF-PJM records the Pennsylvania-New Jersey-Maryland market from 2013-01-01 to 2018-12-24, containing zonal electricity prices in the Commonwealth Edison (COMED) as a target variable, with corresponding system load and COMED load forecasts as known covariates. (3) EPF-BE captures Belgium’s electricity market from 2011-01-09 to 2016-12-31, containing hourly electricity prices as a target variable, with corresponding load forecasts in Belgium and generation forecasts in France as known covariates. (4) EPF-FR documents the electricity market in France from 2012-01-09 to 2017-12-31, containing hourly prices as a target variable, with corresponding load and generation forecasts as known covariates. (5) EPF-DE records the German electricity market from 2012-01-09 to 2017-12-31, containing hourly prices as a target variable, with zonal load forecasts in the TSO Amprion zone and wind and solar generation forecasts as known covariates. (6) The EDF dataset [2] is an hourly time series dataset spanning over five years from 2015-01-03 to 2020-06-27, where the target variable is the national electricity demand in Panama. Observed covariates include four meteorological indices (temperature, relative humidity, liquid precipitation, wind speed) measured in three cities (Tocumen, Santiago, David) in Panama, resulting in a total of 12 indices. In addition, we adopt three binary calendar indicators (holiday, weekend, school day) as known covariates.
In addition, we perform long-term multivariate forecasting on 7 real-world datasets: (1) ETT [44] contains 7 monitoring factors in the electricity transformers from July 2016 to July 2018. ETTh1 and ETTh2 are hourly subsets, and ETTm1 and ETTm2 are subsets recorded every 15 minutes. (2) Weather [40] includes 21 meteorological factors recorded every 10 minutes from the Weather Station of the Max Planck Biogeochemistry Institute in 2020. (3) ECL [40] comprises hourly electricity consumption data from 321 clients. (4) Traffic [40] records hourly road occupancy rates measured by 862 sensors on San Francisco Bay area freeways from January 2015 to December 2016. We use all the variables in these datasets as target variables. The details of datasets are provided in Table 4.
Table 4: Dataset descriptions. #Target, #Observed, and #Known refer to the number of target variables, observed covariates, and known covariates, respectively. The dataset size is presented as (Train, Validation, Test).
Task Dataset #Target #Observed #Known Sampling Frequency Dataset Size
EPF-NP 1 0 2 1 Hour (36500, 5219, 10460)
EPF-PJM 1 0 2 1 Hour (36500, 5219, 10460)
Covariate-Informed EPF-BE 1 0 2 1 Hour (36500, 5219, 10460)
Forecasting EPF-FR 1 0 2 1 Hour (36500, 5219, 10460)
EPF-DE 1 0 2 1 Hour (36500, 5219, 10460)
EDF 1 12 3 1 Hour (33442, 4783, 9586)
ETTh 7 0 0 1 Hour (8545, 2881, 2881)
Multivariate ETTm 7 0 0 15 Minutes (34465, 11521, 11521)
Forecasting ECL 321 0 0 1 Hour (18317, 2633, 5261)
Weather 21 0 0 10 Minutes (36792, 5271, 10540)
Traffic 862 0 0 1 Hour (12185, 1757, 3509)
A.2 Experiment settings
All the experiments are implemented in PyTorch [30]. We utilize a single NVIDIA A100 80GB GPU for the ECL and Traffic datasets and a single NVIDIA V100 32GB GPU for the other datasets. We adopt ADAM [17] with an initial learning rate of $10^{-4}$ and L2 loss for model optimization. The training process is fixed at 10 epochs with early stopping. We set the number of layers in our proposed model $L\in\{1,2,3,4\}$, the number of heads in the multi-head attention layers as 8, dimension of token embedding $D\in\{128,256,512,1024\}$, and a smoothing factor for Attention Score Smoothing $\alpha\in\{0.1,0.2\}$. The patch length is uniformly set to 24 for short-term covariate-informed forecasting and 96 for long-term multivariate forecasting.
A.3 Baseline Descriptions
We cite the baseline results in Table 1 from the TimeXer [34] paper and the results in Tables 2 and 5 from the Timer-XL [24] paper, if available. Otherwise, we reproduced the baseline models based on the TimesNet [39] repository. Note that Timer-XL, TimeXer, and iTransformer are extended from their original implementations to support future known covariates in short-term covariate-informed forecasting. Specifically, for Timer-XL, we modified the masking scheme in TimeAttention so that target token embeddings can attend to known covariate token embeddings from one patch step ahead. For TimeXer and iTransformer, we introduced an additional variate embedding layer for known covariates to handle their longer lengths. On the other hand, Crossformer, TimesNet, and Autoformer remain in their original implementations, which cannot accommodate future covariates as the extension is not straightforward.
Appendix B Full Results
B.1 Full Results of Multivariate forecasting
For a comprehensive comparison, we evaluate our model on well-recognized long-term multivariate forecasting datasets. We adopt the rolling forecasting approach [23], where a model trained with predetermined input and forecasting lengths is evaluated with a longer forecasting length by iteratively integrating its output into the subsequent input. Following previous works [23, 24], we set the input length to 672, the forecasting length to {96, 192, 336, 720}, and the patch length to 96. The full results are listed in Table 5.
Table 5: Full results of the multivariate forecasting, where the input length is set to 672, and the forecasting length is set to {96, 192, 336, 720}. Results are cited from Timer-XL [24] if available; otherwise reproduced. #$1^{\text{st}}$ denotes the first count, and #$1^{\text{st}},2^{\text{nd}}$ denotes the sum of the first and second counts.
Models
CITRAS
Timer-XL
TimeXer
iTrans.
TFT
Cross.
TimesNet
Auto.
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTh1
96
0.355
0.395
0.364 0.397
0.391
0.425
0.387
0.418
0.585
0.542
0.391
0.430
0.452
0.463
0.467
0.499
192
0.386
0.416
0.405 0.424
0.420
0.443
0.416
0.437
0.590
0.546
0.425
0.451
0.474
0.477
0.492
0.523
336
0.403
0.429
0.427 0.439
0.441
0.459
0.434
0.450
0.594
0.551
0.455
0.472
0.493
0.489
0.519
0.531
720
0.412
0.445
0.439 0.459
0.460
0.483
0.447
0.473
0.642
0.586
0.520
0.520
0.560
0.534
0.589
0.560
Avg
0.389
0.421
0.409 0.430
0.428
0.453
0.421
0.445
0.603
0.556
0.448
0.468
0.495
0.491
0.517
0.528
ETTh2
96
0.293
0.358
0.301
0.354
0.281
0.350
0.307
0.365
0.392
0.431
0.533
0.522
0.328
0.386
0.375
0.425
192
0.351 0.395
0.360
0.396
0.339
0.391
0.376
0.406
0.403
0.438
0.707
0.607
0.370
0.410
0.386
0.433
336
0.372
0.416
0.382
0.419
0.371
0.419
0.416
0.435
0.406
0.442
0.884
0.684
0.367
0.417
0.391
0.439
720
0.419
0.453
0.443
0.468
0.422
0.459
0.432
0.456
0.419 0.448
1.132
0.782
0.384
0.436
0.435
0.473
Avg
0.359
0.405
0.372
0.409
0.353
0.405
0.383
0.415
0.405
0.440
0.814
0.649
0.362
0.412
0.397
0.442
ETTm1
96
0.282
0.340
0.297 0.347
0.300
0.358
0.312
0.367
0.491
0.478
0.311
0.365
0.327
0.375
0.517
0.489
192
0.329
0.369
0.344
0.376 0.342
0.381
0.351
0.390
0.505
0.484
0.345
0.386
0.361
0.394
0.523
0.493
336
0.365
0.391
0.382
0.399
0.383
0.402
0.389
0.412
0.592
0.515
0.379
0.407
0.391
0.413
0.551
0.505
720
0.423
0.425
0.450
0.437
0.450
0.436
0.458
0.449
0.576
0.516
0.436
0.440
0.438
0.442
0.608
0.535
Avg
0.350
0.382
0.368 0.390 0.368
0.394
0.377
0.404
0.541
0.498
0.368
0.399
0.379
0.406
0.550
0.505
ETTm2
96
0.174
0.260
0.180
0.263 0.176 0.263
0.183
0.272
0.318
0.360
0.238
0.342
0.207
0.290
0.289
0.363
192
0.231
0.299
0.244
0.306
0.235 0.301
0.241
0.310
0.360
0.383
0.321
0.399
0.267
0.326
0.313
0.377
336
0.280
0.333
0.299
0.344
0.287 0.335
0.296
0.346
0.418
0.414
0.419
0.461
0.331
0.362
0.341
0.394
720
0.365
0.389
0.377
0.397
0.368
0.386
0.393
0.409
0.459
0.440
0.607
0.567
0.406
0.410
0.410
0.438
Avg
0.262
0.320
0.275
0.327
0.266 0.321
0.278
0.334
0.389
0.399
0.396
0.442
0.303
0.347
0.338
0.393
ECL
96
0.128 0.222
0.127
0.219
0.131
0.231
0.133
0.229
0.193
0.294
0.135
0.237
0.184
0.288
0.256
0.357
192
0.149 0.242
0.145
0.236
0.149
0.248
0.158
0.258
0.197
0.298
0.155
0.254
0.192
0.295
0.291
0.376
336
0.164 0.259
0.159
0.252
0.168
0.268
0.168
0.262
0.199
0.301
0.173
0.272
0.200
0.303
0.290
0.379
720
0.199 0.288
0.187
0.277
0.214
0.309
0.205
0.294
0.208
0.309
0.210
0.306
0.228
0.325
0.320
0.403
Avg
0.160 0.253
0.155
0.246
0.166
0.264
0.164
0.258
0.199
0.300
0.168
0.267
0.201
0.303
0.289
0.379
Weather
96
0.152
0.205
0.157
0.205 0.150
0.203
0.174
0.225
0.238
0.288
0.146
0.212
0.169
0.228
0.355
0.409
192
0.200
0.249
0.206
0.250
0.192
0.242
0.227
0.268
0.279
0.316
0.190
0.257
0.222
0.269
0.421
0.450
336
0.253
0.291
0.259
0.291
0.240
0.280
0.290
0.309
0.324
0.343
0.241
0.304
0.290
0.310
0.452
0.465
720
0.325
0.342
0.337
0.344
0.309
0.329
0.374
0.360
0.390
0.384
0.321
0.373
0.376
0.364
0.513
0.496
Avg
0.233
0.272
0.240
0.273
0.223
0.263
0.266
0.291
0.308
0.333
0.225
0.286
0.264
0.293
0.435
0.455
Traffic
96
0.338
0.235
0.340 0.238
0.362
0.264
0.353
0.259
0.574
0.314
0.490
0.282
0.593
0.315
0.675
0.412
192
0.365
0.247
0.360
0.247
0.381
0.275
0.373
0.267
0.583
0.317
0.500
0.287
0.596
0.317
0.679
0.423
336
0.392
0.258
0.377
0.256
0.402
0.288
0.386
0.275
0.591
0.321
0.513
0.293
0.600
0.319
0.688
0.440
720
0.419
0.276
0.418
0.279
0.457
0.324
0.425
0.296
0.614
0.334
0.550
0.314
0.619
0.335
0.693
0.457
Avg
0.379
0.254
0.374
0.255
0.401
0.288
0.384
0.274
0.590
0.322
0.513
0.294
0.602
0.322
0.684
0.433
#$1^{\text{st}}$
16
20
9
7
6
9
0
0
0
0
2
0
2
1
0
0
#$1^{\text{st}},2^{\text{nd}}$
28
33
17 22 17
14
1
0
1
1
8
0
2
2
0
0