### v3

# CITRAS: Covariate-Informed Transformer for Time Series Forecasting

Abstract
In practical time series forecasting, covariates provide rich contextual information that can potentially enhance the forecast of target variables. Although some covariates extend into the future forecasting horizon (e.g., calendar events, discount schedules), most multivariate models fail to leverage this pivotal insight due to the length discrepancy with target variables. Additionally, capturing the dependency between target variables and covariates is non-trivial, as models must precisely reflect the local impact of covariates while also capturing global cross-variate dependencies. To overcome these challenges, we propose CITRAS, a decoder-only Transformer that flexibly leverages multiple targets, past covariates, and future covariates. While preserving strong autoregressive capabilities, CITRAS introduces two novel mechanisms in patch-wise cross-variate attention: Key-Value (KV) Shift and Attention Score Smoothing. KV Shift seamlessly incorporates future covariates into the forecasting of target variables based on their concurrent dependencies. Additionally, Attention Score Smoothing refines locally accurate patch-wise cross-variate dependencies into global variate-level dependencies by smoothing the past series of attention scores. Experimentally, CITRAS outperforms state-of-the-art models on thirteen real-world benchmarks from both covariate-informed and multivariate settings, demonstrating its versatile ability to leverage cross-variate and cross-time dependencies for improved forecasting accuracy.
1 Introduction
Time series forecasting is a cornerstone in diverse fields such as energy (Weron 2014), retail (Böse et al. 2017), and finance (Capistrán, Constandse, and Ramos-Francia 2010), where accurate predictions can drive strategic decisions and enhance operational efficiency. In practical scenarios, forecasters have access not only to the target variables they aim to predict but also to covariates that offer rich contexts for predictions. For instance, the ”Electricity Demand” data shown in Figure 1 (Left) demonstrates a strong negative correlation with the ”Holiday” indicator, which is accessible up to the future forecasting horizon. This highlights the critical importance of appropriately incorporating covariates into the forecasting process.
Driven by these practical demands, this paper addresses time series forecasting that includes two types of covariates. The first is the observed covariate, which has recorded values from the past up to the point of prediction. The second is the known covariate, which has values available from the past through to the future forecasting horizon. Known covariates can take predetermined (e.g., calendar events), estimated (e.g., weather forecasts), or controllable values (e.g., discounts in retail). For clarity, we sometimes refer to the past values of observed covariates and known covariates collectively as past covariates, and the future values of known covariates as future covariates.
Figure 1: Left: Hourly electricity demand data for three weeks in Panama (Aguilar Madrid and Antonio 2021). ”Holiday” and ”Weekend” indicators demonstrate a negative correlation with demand, providing important contextual information in the future forecasting horizon. ”Humidity” displays a weak correlation but introduces temporal disturbances (red circle), which may distract the forecasting process. Right: Performance improvement of CITRAS through the utilization of past covariates and future covariates.
This practical problem setting presents two significant challenges. First, the forecasting model must accommodate heterogeneous variables flexibly. The availability and number of variables can vary depending on the scenario, and known covariates exhibit a length discrepancy with other variables. Second, the model must capture the dependencies between variables from both fine-grained and coarse-grained perspectives. The fine-grained perspective enables the model to precisely capture the local impact of covariates, such as sudden drops in electricity demand on holidays. However, the model that relies solely on a fine-grained perspective may fail to capture global dependencies between variables, as it can overlook covariates representing infrequent events or can be distracted by temporal disturbances. Therefore, it is crucial for the model to capture dependencies between variables at both fine-grained and coarse-grained levels to ensure robust and accurate forecasting.
In recent years, deep forecasting models, particularly Transformers (Vaswani et al. 2017), have made remarkable progress. However, most Transformer-based models have primarily focused on multivariate settings that consider only multiple target variables. Some pioneering works (Wang et al. 2024b; Liu et al. 2025) have highlighted the importance of covariates; yet, the effective integration of heterogeneous covariates that leverages the inherent expressive power of Transformers requires further exploration. Existing approaches to capture cross-variate dependencies can be broadly categorized into two types: variate-level and patch-level. iTransformer (Liu et al. 2024b) is a representative variate-level approach that captures inter-variate correlations by embedding each sequence into a single variate token and applying attention across multiple variate tokens. On the other hand, patch-level approaches, represented by Crossformer (Zhang and Yan 2022), applies cross-variate attention between locally semantic tokens obtained through patching. While variate-level approaches excel at capturing global dependency robustly against noisy interaction, patch-level approaches are better at capturing fine-grained information. Unfortunately, an architecture that enjoys the advantages of both of these approaches is yet to be developed.
To achieve effective time series forecasting that captures the intricate dependencies among diverse variables, we propose CITRAS: Covariate-Informed Transformer for time series forecasting. This model is a decoder-only patch-based Transformer that has a cross-time attention module and a cross-variate attention module separately. To effectively model the cross-variate dependency at both fine-grained and coarse-grained perspectives, we introduce Attention Score Smoothing in the cross-variate attention module, which refines locally accurate patch-wise dependencies into global variate-level dependencies. Furthermore, we introduce Key-Value (KV) Shift mechanism, which associates the key of the known covariate with the value from one patch step ahead, thereby seamlessly integrating future covariates into the prediction process. These innovations expand the covariate understanding capabilities of the canonical decoder-only Transformer, fully preserving its strong autoregressive properties.
In summary, our contributions are as follows:
  • To fully utilize the rich context offered by covariates, we extend the decoder-only Transformer to flexibly accommodate observed covariates and known covariates while fully maintaining its original strong autoregressive capability.
  • Our model, CITRAS, introduces two novel mechanisms into the cross-variate attention module. Attention Score Smoothing captures locally accurate patch-wise dependencies and refines them into global variate-level dependencies, while KV Shift seamlessly integrates future covariates based on the obtained dependencies.
  • Experimentally, CITRAS outperforms state-of-the-art models on thirteen benchmarks from both covariate-informed and multivariate settings, demonstrating its versatile ability to leverage cross-variate and cross-time dependencies for improved forecasting accuracy.
2 Related Work
2.1 Transformer-based Time Series Forecasting
Given the significant success of Transformers in the fields of natural language processing (Devlin et al. 2018) and computer vision (Dosovitskiy et al. 2021), numerous studies have attempted to leverage their capabilities for time series forecasting. These works can be roughly divided into two approaches: channel-dependent (CD) and channel-independent (CI). The CD approach assumes that the future values of a variable are determined by its past values as well as the values of other variables. In contrast, the CI approach posits that a specific variable only depends on its past values, omitting explicit interactions between variables.
Early applications of Transformers in time series forecasting can be viewed as CD approaches, where multiple variables at the same time point are embedded into temporal tokens (Zhou et al. 2022; Liu et al. 2022b; Dong et al. 2023). To address the computational costs and intricate dependencies arising from long time series data, Informer (Zhou et al. 2021) designs a ProbSparse self-attention mechanism with a distilling operation to efficiently focus on important features. Autoformer (Wu et al. 2021) incorporates decomposition and an auto-correlation mechanism to uncover reliable dependencies from complex temporal patterns. However, point-wise temporal tokens have limited local semantics, making it difficult to capture intricate cross-time and cross-variate dependencies.
In response to these challenges, PatchTST (Nie et al. 2023) adopts a CI approach and introduces patching, which handles time series data by segmenting it into fixed-length segments. Patching reduces computational costs and enhances the local semantics in token representation, making it a prevalent technique in subsequent methods (Zhang et al. 2024b, a; Zhou et al. 2024). Additionally, the CI approach is also adopted by many recent large models (Ansari et al. 2024; Das et al. 2024; Goswami et al. 2024) and LLM-based forecasters (Gruver et al. 2023; Zhou et al. 2023; Jin et al. 2024), as it avoids the risk of unnecessary noisy interactions between channels.
However, the CI approach is sometimes regarded as an oversimplification (Abdelmalak et al. 2025), especially considering that fluctuations in time series data are often influenced by external factors. A representative CD approach is iTransformer (Liu et al. 2024b), which embeds each channel into a single variate token and applies cross-variate attention over them. TimeXer (Wang et al. 2024b) and Leddam (Yu et al. 2024) also capture cross-variate dependencies at the variate level, with the additional cross-time attention to capture intra-variate temporal dependencies. Meanwhile, Crossformer (Zhang and Yan 2022) applies cross-variate attention to patch-level representations. Any-variate Attention techniques (Woo et al. 2024; Liu et al. 2025, 2024a) also apply attention to patch-level representations obtained by flattening all variables into one sequence. While the patch-level approach captures more fine-grained dependencies between variables, its local receptive field lacks global context awareness. Our method, CITRAS, first captures fine-grained dependencies at the patch level, and Attention Score Smoothing gradually refines them to achieve global variate-level dependencies.
2.2 Covariate-Informed Time Series Forecasting
To leverage the rich context offered by covariates, numerous approaches have been proposed, including statistical models (Williams 2001; Nicholson, Matteson, and Bien 2017), MLP variants (Das et al. 2023; Olivares et al. 2023; Chen et al. 2023; Ekambaram et al. 2024), RNN variants (Salinas et al. 2020), and Transformer-based models (Lim et al. 2021). For instance, TiDE (Das et al. 2023) introduces an MLP-based architecture that utilizes future covariates in the final projection to leverage their direct effects on future target variables. TSMixer-Ext (Chen et al. 2023) embeds past and future covariates into the same shape and applies feature mixing in the early stage. Temporal Fusion Transformer (Lim et al. 2021) designs gating mechanisms on top of the Transformer to select relevant features for forecasting. ChronosX (Arango et al. 2025) introduces injection blocks to integrate covariates into pretrained models. A common drawback of these models is their structural complexity, which arises from architectural additions needed to handle heterogeneous variables.
Recent works, such as TimeXer (Wang et al. 2024b) and Timer-XL (Liu et al. 2025), have attempted simpler integration of covariates into Transformers to retain their inherent expressive power. However, these approaches require further exploration. Specifically, the variate-level approach adopted by TimeXer is not compatible with decoder-only Transformers, leaving room for enhanced autoregressive capabilities through next-token prediction. The masking scheme in Timer-XL, known as TimeAttention, is not designed to accommodate future covariates, raising questions about its ability to utilize these pivotal features. In contrast to these works, CITRAS seamlessly integrates both observed and known covariates into the decoder-only Transformer, fully preserving its strong autoregressive capability.
3 CITRAS
3.1 Problem Setting
Let $\mathbf{X}_{1:T}^{tgt,:}=\{\mathbf{X}_{1:T}^{tgt,1},\mathbf{X}_{1:T}^{tgt,2},...,\mathbf{X}_{1:T}^{tgt,C_{tgt}}\}\in\mathbb{R}^{T\times C_{tgt}}$ be a target time series of length $T$ and a number of variables $C_{tgt}$. With optional usage of observed covariates $\mathbf{X}_{1:T}^{obs,;}\in\mathbb{R}^{T\times C_{obs}}$ and known covariates $\mathbf{X}_{1:T+S}^{knw,:}\in\mathbb{R}^{(T+S)\times C_{knw}}$, the goal of the forecasting model $\mathcal{F}_{\theta}$ is to predict the future $S$ time steps of the target time series, which can be formulated as follows:
$\widehat{\mathbf{X}}_{T+1:T+S}^{tgt,:}=\mathcal{F}_{\theta}\left(\mathbf{X}_{1:T}^{tgt,:},\mathbf{X}_{1:T}^{obs,:},\mathbf{X}_{1:T+S}^{knw,:}\right)$ (1)
3.2 Architecture
As shown in Figure 2, CITRAS is a patch-based, decoder-only Transformer with $L$ layers of cross-time attention modules and cross-variate attention modules. Our core mechanisms, KV Shift and Attention Score Smoothing, reside in the cross-variate attention module.
Figure 2: Overall structure of CITRAS. (a) The embedding module applies patching to each variable, yielding temporal token embeddings. (b) The cross-time attention module captures intra-variate cross-time dependencies. (c) The cross-variate attention module captures cross-variate dependencies. (d) KV Shift associates the key of the known covariate with the value from one patch step ahead. Attention Score Smoothing calculates the Exponential Moving Average (EMA) of patch-wise attention scores. (e) The projection layer projects each target token embedding to the values of the next patch.
Embedding
First, all sequences are segmented into non-overlapping patches of length $P$, and each patch is embedded into a $D$-dimensional token. For simplicity, we assume that both $T$ and $S$ are divisible by $P$. Taking target variables $c$ as an example, this can be formalized as:
$\{\mathbf{s}_{1}^{tgt,c},\mathbf{s}_{2}^{tgt,c},.,\mathbf{s}_{N_{tgt}}^{tgt,c}\}=\operatorname{Patchfy}\left(\mathbf{X}_{1:T}^{tgt,c}\right)$ (2)
$\mathbf{H}_{i}^{tgt,c}=\operatorname{Embed}\left(\mathbf{s}_{i}^{tgt,c}\right),~~i=1,.,N_{tgt}$
where $N_{tgt}=\frac{T}{P}$ is the number of patches and $\operatorname{Embed}:\mathbb{R}^{P}\to\mathbb{R}^{D}$ is a shared linear projector across all variables. We denote the token embeddings of a target variable at all patch steps as $\mathbf{H}_{:}^{tgt,c}=~\{\mathbf{H}_{i}^{tgt,c}\}_{i=1}^{N_{tgt}}~\in~\mathbb{R}^{N_{tgt}\times D}$. Similarly, the token embeddings of an observed covariate are denoted as $\mathbf{H}_{:}^{obs,c}\in\mathbb{R}^{N_{obs}\times D}$, and those of a known covariate are denoted as $\mathbf{H}_{:}^{knw,c}\in\mathbb{R}^{N_{knw}\times D}$, all obtained using the shared parameters with the target variables. Here, $N_{obs}=N_{tgt}=\frac{T}{P}$ and $N_{knw}=\frac{(T+S)}{P}$.
Cross-Time Attention
In the cross-time attention module, we apply multi-head attention with causal masking to all variables to capture their intra-variate cross-time dependency. Following TimerXL (Liu et al. 2025), we adopt Rotary Position Embedding (RoPE) (Su et al. 2024) to capture temporal order. Taking target variables $c$ as an example, and dropping layer index for brevity, this can be formalized as:
$\widetilde{\mathbf{H}}_{:}^{tgt,c}$ $=\operatorname{LN}\left(\mathbf{H}_{:}^{tgt,c}+\operatorname{MHA}\left(\mathbf{H}_{:}^{tgt,c},\mathbf{H}_{:}^{tgt,c},\mathbf{H}_{:}^{tgt,c}\right)\right)$ (3)
$\mathbf{H}_{:}^{tgt,c}$ $=\operatorname{LN}\left(\widetilde{\mathbf{H}}_{:}^{tgt,c}+\operatorname{FFN}\left(\widetilde{\mathbf{H}}_{:}^{tgt,c}\right)\right)$
where $\operatorname{LN}$ denotes layer normalization (Ba, Kiros, and Hinton 2016), $\operatorname{MHA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)$ denotes the multi-head attention layer where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ serve as queries, keys and values, and $\operatorname{FFN}$ denotes a feed-forward network.
Again, token embeddings of observed covariates $\mathbf{H}_{:}^{obs,c}$ and those of known covariates $\mathbf{H}_{:}^{knw,c}$ are similarly processed with the shared parameters.
Cross-Variate Attention — KV Shift
Attention is recognized as an effective approach for capturing cross-variate dependencies (Liu et al. 2024b). To represent the dependencies of target variables while avoiding unnecessary interactions among covariates, we adopt multi-head attention, where queries originate from the target variables, whereas keys and values are derived from all available variables. However, applying attention across the variables of the same patch steps, as adopted by Crossformer (Zhang and Yan 2022), fails to leverage pivotal future patches of known covariates, as they do not have corresponding target patches.
To address this issue, we introduce KV Shift mechanism in the multi-head attention layer of the cross-variate attention module, which associates the key of the known covariate with the value from one patch step ahead. Specifically, the key and value at patch step $i$ can be formalized as:
$\text{Key:}~~~\mathbf{H}_{i}^{k,:}$ $=\left[\mathbf{H}_{i}^{tgt,:},\mathbf{H}_{i}^{obs,:},\mathbf{H}_{i}^{knw,:}\right]$ (4)
$\text{Value:}~~~\mathbf{H}_{i}^{v,:}$ $=\left[\mathbf{H}_{i}^{tgt,:},\mathbf{H}_{i}^{obs,:},\mathbf{H}_{i+1}^{knw,:}\right]$
where $\left[\cdot,\cdot\right]$ denotes the concatenation along the variate dimension and $\mathbf{H}_{i}^{k,:},\mathbf{H}_{i}^{v,:}\in\mathbb{R}^{\left(C_{tgt}+C_{obs}+C_{knw}\right)\times D}$. After this, cross-variate dependency can be captured by the standard multi-head attention layer:
$\widetilde{\mathbf{H}}_{i}^{tgt,:}$ $=\operatorname{LN}\left(\mathbf{H}_{i}^{tgt,:}+\operatorname{MHA}\left(\mathbf{H}_{i}^{tgt,:},\mathbf{H}_{i}^{k,:},\mathbf{H}_{i}^{v,:}\right)\right)$ (5)
$\mathbf{H}_{i}^{tgt,:}$ $=\operatorname{LN}\left(\widetilde{\mathbf{H}}_{i}^{tgt,:}+\operatorname{FFN}\left(\widetilde{\mathbf{H}}_{i}^{tgt,:}\right)\right)$
for $i=1,...,N_{tgt}$. KV Shift maintains step alignment in the dot product calculation between queries and keys, enabling precise capture of concurrent dependencies. It then allows the target token to incorporate the next-step future known covariate (value) based on the strength of the obtained dependency. As the target token gradually transforms into the prediction for the next step, this facilitates a natural flow of information for leveraging future known covariates.
Cross-Variate Attention — Attention Score Smoothing
The attention score of $\mathrm{MHA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)$ is calculated as:
$\widetilde{\mathbf{A}}=\left(\mathbf{Q}\mathbf{W}_{q}\right)\left(\mathbf{K}\mathbf{W}_{k}\right)^{\top}$ (6)
where $\mathbf{W}_{q},\mathbf{W}_{k}\in\mathbb{R}^{D\times d_{k}}$ and $d_{k}$ is the dimension of the query, key, and value. In the cross-variate attention module, this attention score is calculated for each patch step separately, capturing fine-grained dependency between variables at that step. However, time series data often exhibit local disturbances, and sparse variables representing infrequent events may maintain constant values within a patch. In such scenarios, local patch-wise attention fails to represent global variate-level dependencies.
To address this issue, we introduce Attention Score Smoothing in the multi-head attention layer of the cross-variate attention module, which transforms locally accurate patch-wise dependencies into global variate-level dependencies. Specifically, the attention score at each step is smoothed based on the attention scores up to that step in each head. We employ an Exponential Moving Average (EMA) as a smoothing method, as it enables adaptation to shifting correlations between variables (Kang, Shin, and Lee 2025) by assigning exponentially decreasing weights over time. Denoting original attention score at patch step $i$ as $\widetilde{\mathbf{A}}_{i}$, the smoothed attention score $\mathbf{A}_{i}$ can be calculated as:
$\mathbf{A}_{i}$ $=\alpha\widetilde{\mathbf{A}}_{i}+\left(1-\alpha\right)\mathbf{A}_{i-1},~~i=2,.,N_{tgt}$ (7)
where $\mathbf{A}_{1}=\widetilde{\mathbf{A}}_{1}$ and $\alpha$ is a smoothing factor, which we consider as a shared hyperparameter across all cross-variate attention modules and heads. It is important to note that the functionality provided by Attention Score Smoothing cannot be achieved by existing approaches that integrate EMA into attention mechanisms, as attempted in ETSformer (Woo et al. 2022) or CARD (Wang et al. 2024a). These works apply temporal smoothing in cross-time attention (i.e., along the parallel direction) to help capture temporal dependencies. In contrast, we apply temporal smoothing in cross-variate attention (i.e., along the orthogonal direction) to refine cross-variate dependencies over time.
Projection
Following the next-token prediction approach common in decoder-only Transformers, each token embedding of the target variables is used to predict the values of the next patch. For $i=1,...,N_{tgt}$ and $c=1,...,C_{tgt}$, this can be formalized as:
$\widehat{\mathbf{X}}_{iP+1:(i+1)P}^{tgt,c}=\operatorname{Project}\left(\mathbf{H}_{i}^{tgt,c}\right)$ (8)
where $\operatorname{Project}:\mathbb{R}^{D}\to\mathbb{R}^{P}$ is a shared linear projector across all steps and target variables. In the training phase, we calculate the squared loss using all of these outputs. In the testing phase, the outputs from the last patches are used for forecasting. When the forecasting horizon $S$ exceeds the patch length $P$, the output target values are integrated into subsequent input for recursive forecasting.
4 Experiments
To verify the effectiveness and versatility of CITRAS, we extensively evaluated it in the two settings, covariate-informed forecasting and well-acknowledged multivariate forecasting.
4.1 Covariate-Informed Forecasting
Dataset
We use seven real world datasets for covariate-informed forecasting. (1)-(5) EPF (Lago et al. 2021): Electricity Price Forecasting datasets of five major power markets spanning six years. The target variable is the hourly electricity price and we adopt day-ahead load forecasts and generation forecasts as known covariates. (6) EDF (Aguilar Madrid and Antonio 2021): Electricity Demand Forecasting dataset in Panama spanning over five years. The target variable is the hourly national electricity demand in Panama. We adopt four meteorological indicators measured in three cities (total 12 indices) as observed covariates, and three binary calendar features (holiday, weekend, school day) as known covariates. (7) BS (Fanaee-T 2013): Bike Sharing dataset recording the hourly rental counts of Bikeshare system in Washington D.C. spanning two years. The target variables are three types of rental counts. We adopt five meteorological indicators as observed covariates and three calendar features (holiday, weekday, working day) as known covariates. Following previous works (Olivares et al. 2023), we set the input length as 168 and forecasting length as 24. Further details about the datasets are included in Appendix A.1.
Baselines
We include the following baselines, categorized into four groups. (A) CD models that originally support future covariates, including TSMixer-Ext (Chen et al. 2023), TiDE (Das et al. 2023), Temporal Fusion Transformer (TFT) (Lim et al. 2021), and TimeXer (Wang et al. 2024b). (B) CD models that are extended from original designs (as detailed in A.3) to support future covariates, including Timer-XL (Liu et al. 2025), iTransformer (Liu et al. 2024b), and Leddam (Yu et al. 2024). (C) CD models that cannot accommodate future covariates, including CARD (Wang et al. 2024a), ModernTCN (Donghao and Xue 2024), TimesNet (Wu et al. 2023), and Crossformer (Zhang and Yan 2022). (D) CI models including FITS (Xu, Zeng, and Xu 2024), DLinear (Zeng et al. 2023), and PatchTST (Nie et al. 2023). For fair comparison with baselines (C) and (D), we also evaluate CITRAS without future covariates (i.e., without using KV Shift).
Result
Table 1 presents the results of covariate-informed forecasting with and without the use of future covariates. In the former setting, CITRAS outperforms TSMixer-Ext, TiDE, and TFT, all of which employ complex architectures to handle future covariates. This superior performance can be attributed to the simple yet effective design of CITRAS, which seamlessly integrates future covariates into the decoder-only Transformer through KV Shift, thereby preserving its inherent expressive power. Among recent models that adhere to the canonical Transformer architecture, TimeXer captures cross-variate dependencies at the variate level, while Timer-XL does so at the patch level. The relative superiority of these contrasting methods is dataset-dependent. In contrast, CITRAS leverages the advantages of both approaches through Attention Score Smoothing, consistently demonstrating strong performance across datasets. Moreover, the superiority of CITRAS remains evident even in conventional settings without future covariates. This is because standard multivariate models suffer from unnecessary interactions among covariates, as they treat target variables and covariates equally. In contrast, CITRAS applies attention only from target variables to covariates, thereby avoiding such negative interactions.
Model
CITRAS
TSMixer-Ext
TiDE
TFT
TimeXer
Timer-XL
iTrans.
Leddam
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
w/ future
EPF–NP
0.171
0.214
0.187
0.253
0.296
0.309
0.193
0.237
0.175
0.225
0.176
0.220
0.203
0.250
0.227
0.266
EPF-PJM
0.063
0.153
0.078
0.174
0.105
0.213
0.079
0.174
0.080
0.175
0.062
0.155
0.077
0.171
0.087
0.193
EPF–BE
0.362
0.248
0.316
0.241
0.447
0.301
0.352
0.240
0.368
0.231
0.384
0.257
0.339
0.246
0.369
0.253
EPF–FR
0.363
0.176
0.396
0.214
0.407
0.251
0.371
0.182 0.364
0.193
0.430
0.214
0.378
0.205
0.416
0.221
EPF–DE
0.217
0.293
0.250
0.322
0.515
0.462
0.262
0.326
0.283
0.322
0.228 0.299
0.258
0.327
0.273
0.332
EDF
0.070
0.193
0.077
0.209
0.133
0.259
0.092
0.225
0.071
0.199
0.079
0.206
0.071 0.197
0.080
0.206
BS
0.280
0.313
0.354
0.363
0.488
0.446
0.318
0.346
0.338
0.357
0.286 0.318
0.358
0.360
0.376
0.384
Model
CITRAS
CARD
ModernTCN
TimesNet
Cross.
FITS
DLinear
PatchTST
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
w/o future
EPF–NP
0.226
0.259
0.256
0.287
0.234 0.272
0.250
0.289
0.240
0.285
0.304
0.316
0.309
0.321
0.267
0.284
EPF-PJM
0.088
0.181
0.108
0.211
0.101
0.195 0.097 0.195
0.101
0.199
0.109
0.216
0.108
0.215
0.106
0.209
EPF–BE
0.405
0.275
0.414
0.278
0.383
0.260
0.419
0.288
0.420
0.290
0.460
0.308
0.463
0.313
0.400 0.262
EPF–FR
0.408
0.220
0.420
0.243
0.407
0.224
0.431
0.234
0.434
0.208
0.402
0.254
0.429
0.260
0.411
0.220
EPF–DE
0.408
0.400
0.453
0.424
0.444 0.416
0.502
0.446
0.574
0.430
0.533
0.473
0.520
0.463
0.461
0.432
EDF
0.086
0.205
0.088
0.211
0.086
0.214
0.101
0.233
0.108
0.233
0.134
0.258
0.129
0.252
0.103
0.232
BS
0.289
0.320
0.416
0.414
0.378
0.402
0.363
0.366
0.327
0.354
0.484
0.439
0.469
0.429
0.378
0.348
Table 1: Full results of the covariate-informed forecasting with and without future covariates. The input length and forecasting length is set to 168 and 24, respectively. EPF results for ”w/o future” are cited from TimeXer  (Wang et al. 2024b) if available; other results are reproduced. The best result is represented in bold, followed by underline.
Model
CITRAS
TimeXer
Timer-XL
iTrans.
Leddam
CARD
ModernTCN
FITS
DLinear
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTh1
0.389
0.421
0.428
0.453
0.409
0.430
0.421
0.445
0.406
0.431
0.414
0.433
0.413
0.432
0.415
0.426
0.426
0.444
ETTh2
0.359
0.405
0.353
0.405
0.372
0.409
0.383
0.415
0.339
0.391
0.340
0.388
0.336
0.388 0.339
0.386
0.352
0.404
ETTm1
0.350
0.382
0.368
0.394
0.368
0.390
0.377
0.404
0.358
0.384
0.355
0.381
0.369
0.393
0.358
0.379
0.357
0.380
ETTm2
0.262
0.320
0.266
0.321
0.275
0.327
0.278
0.334
0.254
0.315
0.247
0.308
0.267
0.326
0.254 0.314
0.262
0.326
ECL
0.160 0.253
0.166
0.264
0.155
0.246
0.164
0.258
0.161
0.253 0.160
0.255
0.183
0.291
0.171
0.269
0.165
0.265
Weather
0.233
0.272
0.223
0.263
0.240
0.273
0.266
0.291
0.225
0.265
0.221
0.266
0.232
0.273
0.229
0.265
0.239
0.291
Traffic
0.379
0.254
0.401
0.288
0.374
0.255
0.384
0.274
0.386
0.270
0.400
0.280
0.446
0.327
0.415
0.286
0.423
0.298
PEMS04
0.126
0.220
0.147
0.261
0.126
0.227
0.135
0.237
0.126
0.228
0.186
0.285
0.141
0.261
0.262
0.340
0.244
0.333
PEMS08
0.236
0.216
0.270
0.279
0.247 0.234
0.280
0.253
0.265
0.244
0.330
0.321
0.337
0.324
0.442
0.378
0.408
0.380
Table 2: Results of the multivariate forecasting, where the input length is set to 672, and the forecasting length is set to {96, 192, 336, 720}. We adopt the rolling forecasting approach  (Liu et al. 2024c), where one model is used for four forecasting lengths. Averaged results are reported here and full results are provided in Table 8. Results are cited from Timer-XL (Liu et al. 2025) if available; otherwise reproduced.
4.2 Multivariate Forecasting
For a comprehensive comparison, we further evaluate our model on well-recognized multivariate forecasting datasets without using any covariates. To reflect practical scenarios, we adopt the rolling forecasting approach (Liu et al. 2024c), where a model trained with predetermined input and forecasting lengths is evaluated with a longer forecasting length by iteratively integrating its output into the subsequent input. Following previous works (Liu et al. 2024c, 2025), we set the input length to 672, the training forecasting length to 96, and the testing forecasting length to {96, 192, 336, 720}. We present the average results for the four forecasting lengths in Table 2. Compared to recent models specifically designed for these multivariate settings, CITRAS achieves superior or competitive performances in most datasets, highlighting the effectiveness and versatility of its design across various settings.
4.3 Model Analysis
Qualitative Analysis
To demonstrate the strong covariate understanding capability of CITRAS, we visually compare the forecasting results without and with the use of covariates, as shown in Fig. 3. In the EPF-NP result, CITRAS accurately predicts the sudden rise in target electricity prices by taking into account the high grid load and low power generation forecasts during the forecasting horizon. In the BS result, it effectively captures the negative impact of the binary working day covariate and accurately forecasts the increase in rental counts on the day following a working day. These results indicate that CITRAS does not simply incorporate covariates that resemble the target shape, but rather captures complex relationships from a global perspective and accurately reflects the local impact of future covariates. This multi-grained utilization of covariates gives CITRAS a distinct advantage over TimeXer, which leverages cross-variate dependencies only at the variate level.
Figure 3: Forecasting examples of CITRAS and TimeXer without (shown in green) and with (shown in blue) the use of covariates. The target variable is displayed in the top box, while the known covariates are shown in the boxes below.
Ablation Study
CITRAS introduces KV Shift and Attention Score Smoothing (ASS) on top of a decoder-only Transformer. To validate the effectiveness of these design choices, we conduct detailed ablation studies, as summarized in Table 3. In the KV Shift ablation, the model fails to incorporate future covariates, leading to a significant performance drop in covariate-informed forecasting. In the ASS ablation, the model is restricted to capturing cross-variate dependencies only from a local perspective, resulting in inferior performance in both covariate-informed and multivariate settings. A more in-depth analysis of the sensitivity to the smoothing factor is provided in Appendix B.1. For the decoder-only ablation, we alternatively adopt an encoder-only architecture (i.e., removing causal masking in cross-time attention and applying a projection layer to the flattened features of all tokens for prediction). However, this design cannot effectively model causality in forecasting and lacks token-wise supervision, thus failing to fully leverage the inherent autoregressive capability of the Transformer in most cases. Collectively, these results highlight the strength of CITRAS, which seamlessly integrates heterogeneous covariates into a decoder-only Transformer and effectively models both cross-variate and cross-time dependencies.
Design
EPF-NP
EDF
BS
ETTh1
ECL
CITRAS
0.171
0.070
0.280
0.389
0.160
(w/o) KV Shift
0.226
0.086
0.289
(w/o) ASS
0.175
0.074
0.299
0.401
0.165
(w/o) Decoder-only
0.188
0.068
0.366
0.428
0.162
Table 3: MSE results of the ablation study.
Representation Analysis
To uncover the effects of Attention Score Smoothing further, we present the visualization of the cross-variate attention scores between the target variable and other variables in the EDF dataset in Figure 4. It is observed that patch-wise attention scores effectively capture the local negative impact of ”Holiday” on the target variable (”Electricity Demand”) . However, these scores are also susceptible to local disturbances, as demonstrated by the relatively fluctuating scores of ”Humidity” compared to the ”Temperature” that shows stable correlation with the target variable. By applying Attention Score Smoothing, the obtained scores effectively convey the strong dependency on ”Holiday” from past steps and mitigate the noisy interaction caused by the fluctuating scores of ”Humidity”. In this way, Attention Score Smoothing enjoys the advantages of both patch-level and variate-level approaches while offering improved interpretability.
Figure 4: Visualization of attention scores between the target ”Electricity Demand” and other variables observed in EDF dataset. By applying Attention Score Smoothing, the smoothened scores capture strong dependency on ”Holiday” while mitigating noisy interaction between ”Humidity”.
5 Conclusion
Considering the critical roles of covariates in practical time series forecasting, we propose CITRAS, a decoder-only Transformer that flexibly leverages multiple targets, observed covariates, and known covariates. Specifically, we introduce two novel mechanisms within the patch-wise cross-variate attention modules: Key-Value Shift, which seamlessly integrates future values of known covariates in to forecasting, and Attention Score Smoothing, which refines locally accurate dependencies into global variate-level dependencies. Experimentally, CITRAS outperforms state-of-the-art models on thirteen benchmarks from both covariate-informed and multivariate settings, demonstrating its versatile ability to leverage cross-variate and cross-time dependencies for improved forecasting accuracy. In the future, we will explore the generalization capability of CITRAS through large-scale pre-training to take advantage of its simple decoder-only architecture.
Appendix A Implementation Details
A.1 Dataset Descriptions
We perform covariate-informed forecasting on seven real world datasets. They include target variable(s), observed covariates, and known covariates, thereby providing a realistic representation of practical forecasting scenarios. EPF datasets (Lago et al. 2021) consist of five subsets from different day-ahead electricity markets, each spanning six years: (1) EPF-NP documents the Nord Pool electricity market from 2013-01-01 to 2018-12-24, containing hourly electricity prices as a target variable, with corresponding grid load and wind power forecasts as known covariates. (2) EPF-PJM records the Pennsylvania-New Jersey-Maryland market from 2013-01-01 to 2018-12-24, containing zonal electricity prices in the Commonwealth Edison (COMED) as a target variable, with corresponding system load and COMED load forecasts as known covariates. (3) EPF-BE captures Belgium’s electricity market from 2011-01-09 to 2016-12-31, containing hourly electricity prices as a target variable, with corresponding load forecasts in Belgium and generation forecasts in France as known covariates. (4) EPF-FR documents the electricity market in France from 2012-01-09 to 2017-12-31, containing hourly prices as a target variable, with corresponding load and generation forecasts as known covariates. (5) EPF-DE records the German electricity market from 2012-01-09 to 2017-12-31, containing hourly prices as a target variable, with zonal load forecasts in the TSO Amprion zone and wind and solar generation forecasts as known covariates. (6) EDF dataset (Aguilar Madrid and Antonio 2021) is an hourly time series dataset spanning over five years from 2015-01-03 to 2020-06-27, where the target variable is the national electricity demand in Panama. Observed covariates include four meteorological indicators (T2M=Air temperature at 2 m ($^{\circ}\mathrm{C}$), QV2M=Specific humidity at 2 m (%), TQL=Total precipitable liquid water ($1/m^{2}$), and W2M=Wind speeed at 2 m ($m/s$)) measured in three cities (Tocumen, Santiago, David) in Panama, resulting in a total of 12 indices. In addition, we adopt three binary calendar features (holiday, school day, weekend) as known covariates. (7) BS (Fanaee-T 2013) is an hourly time series data spanning two years from 2011-01-01 to 2012-12-31, where the target variables are three types of rental counts (casual, registered, and cnt=total count) in Bike Sharing system in Washington D.C. Observed covariates include five meteorological indicators (weathersit=weather situation, temp=temperature, atemp=feeling temperature, hum=humidity, and windspeed). In addition, we adopt three calendar features (holiday, weekday, working day) as known covariates.
In addition, we perform multivariate forecasting on nine real-world datasets: ETT (Zhou et al. 2021) contains 7 monitoring factors in the electricity transformers from July 2016 to July 2018. (1) ETTh1 and (2) ETTh2 are hourly subsets, and (3) ETTm1 and (4) ETTm2 are subsets recorded every 15 minutes. (5) Weather (Wu et al. 2021) includes 21 meteorological factors recorded every 10 minutes from the Weather Station of the Max Planck Biogeochemistry Institute in 2020. (6) ECL (Wu et al. 2021) comprises hourly electricity consumption data from 321 clients. (7) Traffic (Wu et al. 2021) records hourly road occupancy rates measured by 862 sensors on San Francisco Bay area freeways from January 2015 to December 2016. (8) PEMS04 (Liu et al. 2022a) contains the public traffic network data recorded at 307 locations in California every 5 minutes from 2018-01-01 to 2018-02-28. Similarly, (9) PEMS08 (Liu et al. 2022a) contains the data at 170 locations from 2016-07-01 to 2016-08-31. We use all the variables in these datasets as target variables. The summary of datasets are provided in Table 4.
Task Dataset #Target #Observed #Known Sampling Frequency Dataset Size
Covariate-informed Forecasting EPFs 1 0 2 1 Hour (36500, 5219, 10460)
EDF 1 12 3 1 Hour (33442, 4783, 9586)
BS 3 5 3 1 Hour (11974, 1716, 3452)
Multivariate Forecasting ETTh 7 0 0 1 Hour (8545, 2881, 2881)
ETTm 7 0 0 15 Minutes (34465, 11521, 11521)
ECL 321 0 0 1 Hour (18317, 2633, 5261)
Weather 21 0 0 10 Minutes (36792, 5271, 10540)
Traffic 862 0 0 1 Hour (12185, 1757, 3509)
PEMS04 307 0 0 5 Minutes (10100, 3400, 3399)
PEMS08 170 0 0 5 Minutes (10618, 3573, 3572)
Table 4: Dataset descriptions. #Target, #Observed, and #Known refer to the number of target variables, observed covariates, and known covariates, respectively. The dataset size is presented as (Train, Validation, Test).
A.2 Experiment Settings
All the experiments are implemented in PyTorch (Paszke et al. 2019). We utilize a single NVIDIA A100 80GB GPU for the ECL, Traffic, and PEMS datasets and a single NVIDIA V100 32GB GPU for the other datasets. We apply Series Stationarization (Liu et al. 2022c) to the embedding and projection layers of CITRAS to mitigate nonstationarity. We adopt ADAM (Kingma and Ba 2014) with an initial learning rate of $10^{-4}$ and L2 loss for model optimization. The training process is fixed at 10 epochs with an early stopping mechanism, which terminates training if the validation performance does not improve for 3 consecutive epochs. We set the number of layers in our proposed model $L\in\{1,2,3,4\}$, the number of heads in the multi-head attention layers as 8, dimension of token embedding $D\in\{128,256,512,1024\}$, and the smoothing factor for Attention Score Smoothing $\alpha\in\{0.1,0.2,0.4,0.8\}$. The patch length is uniformly set to 24 for covariate-informed forecasting and 96 for multivariate forecasting. We ensure that the drop last option is set to False in the test data loader to avoid any unfair impact, as pointed out by Qiu et al.. All evaluations are conducted with a fixed random seed and a single run for each experiment.
A.3 Baseline Descriptions
We reproduced the baseline models based on the TimesNet (Wu et al. 2023) repository or official repositories. Note that Timer-XL, TimeXer  [^1] The TimeXer implementation that utilizes future covariates is not officially released., iTransformer, and Leddam are extended from their original implementations to support future covariates in covariate-informed forecasting. Specifically, for Timer-XL, we modified the masking scheme in TimeAttention so that target token embeddings can attend to known covariate token embeddings from one patch step ahead. For TimeXer, iTransformer, and Leddam, we introduced an additional variate embedding layer for known covariates to handle their longer lengths.
Appendix B Extra Model Analysis
B.1 Attention Score Smoothing
Attention Score Smoothing captures locally accurate patch-level cross-variate dependency and refines it into global variate-level dependency. Note that a large smoothing factor alpha emphasizes patch-level dependency (with $\alpha=1.0$ corresponding to no smoothing), while a smaller alpha emphasizes variate-level dependency through heavier smoothing. Figure 5 illustrates the performance at different levels of granularity in cross-variate attention. The performance of the variate-level approach is represented by TimeXer. We observe that the suitability of variate-level versus patch-level approaches is dependent on the dataset. However, strong Attention Score Smoothing (around $\alpha=0.2$) strikes a balance between the two, highlighting the versatile effectiveness of this mechanism in both covariate-informed and multivariate forecasting.
Note that this smoothing factor can be treated as a learnable parameter. However, as shown in Table 5, we empirically find that a fixed smoothing factor is sufficient to achieve the desired smoothing effect. Consequently, our Attention Score Smoothing introduces no new learning parameters.
Figure 5: The performance at different levels of granularity in cross-variate attention. The performance at ”Variate” level is represented by TimeXer. Strong Attention Score Smoothing (around $\alpha=0.2$) outperforms both variate-level and patch-level approach.
Smoothing factor
EDF
BS
ETTh1
ECL
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Learnable
0.070
0.193
0.281
0.315
0.399
0.427
0.178
0.270
Fixed
0.070
0.193
0.280
0.313
0.389
0.421
0.160
0.253
Table 5: Comparison between learnable and fixed smoothing factor.
B.2 Computational Efficiency
Consider the covariate-informed setting where the number of target variables is $1$, the number of covariates (including observed and known covariates) is $C$, the patch length of the target variable is $N$. CITRAS calculates cross-variate attention between the target and covariates at each patch step, resulting in a computational complexity of $O(CN)$. Additionally, it applies cross-time attention to each variate separately, incurring a complexity of $O(CN^{2})$. This is favorable compared to the Any-variate Attention adopted by Timer-XL, which incurs $O(C^{2}N^{2})$ to flexibly accommodate covariates. Furthermore, the $O(CN)$ complexity of cross-variate attention is advantageous when the number of covariates $C$ increases, compared to iTransformer, which increases to $O(C^{2})$. We acknowledge that CITRAS has quadratic complexity in the time dimension. Also, the complexity of cross-variate attention can be quadratic with respect to the number of target variables. However, we can introduce advanced attention modules, such as linear attention, to reduce the complexity, as it is based on the simple Transformer architecture.
Besides the theoretical analysis, we demonstrate the empirical model efficiency on the BS dataset and PEMS08 dataset in Fig. 6. We use the best model configuration for each model and evaluate using the same batch size for a fair comparison. The results in the BS dataset underscore the efficient model design of CITRAS. In the PEMS08 dataset, we observe the trade-off between training time and model performance in terms of MSE. Among them, CITRAS achieves the best performance without significant compromise in efficiency.
Figure 6: Comparison of model efficiency on the BS and PEMS08 datasets. Note that the x-axis is in logarithmic scale for the PEMS08 result.
Appendix C Extra Results
C.1 Error Bar
We report the standard deviation of CITRAS performance on covariate-informed forecasting with future known covariates under five runs with different random seeds. The result in Table 6 exhibits that the performance of CITRAS is stable.
Model
CITRAS
Metric
MSE
MAE
EPF–NP
0.171
$\pm$0.004
0.214
$\pm$0.004
EPF-PJM
0.0625
$\pm$0.0013
0.153
$\pm$0.001
EPF–BE
0.362
$\pm$0.010
0.248
$\pm$0.004
EPF–FR
0.363
$\pm$0.010
0.176
$\pm$0.002
EPF–DE
0.217
$\pm$0.003
0.293
$\pm$0.001
EDF
0.0704
$\pm$0.0005
0.193
$\pm$0.001
BS
0.280
$\pm$0.001
0.313
$\pm$0.001
Table 6: Error bar of CITRAS on the covariate-informed forecasting with future covariates.
C.2 Comparison with Pre-trained Models
We further compare CITRAS with large pre-trained models under covariate-informed settings with future covariates. For baselines, we choose Moirai (Woo et al. 2024) (moirai-1.1-R-large) and Moirai-MoE (Liu et al. 2024a) (moirai-moe-1.0-R-base), which are state-of-the-art pre-trained models that natively support future covariates. For a fair comparison, we present the zero-shot results of these models and the full-shot results of Moirai, which is fine-tuned with the full training dataset for a maximum of five epochs with a learning rate $lr\in\{1e-5,1e-6,1e-7\}$. Table 7 shows that CITRAS consistently outperforms Moirai under full-shot settings. Note that the implementation of ChronosX (Arango et al. 2025), which fine-tunes univariate pre-trained models to integrate covariates, is not publicly available yet and therefore not comparable.
Setting
Model
EPF-NP
EDF
BS
MSE
MAE
MSE
MAE
MSE
MAE
full
CITRAS
0.171
0.214
0.070
0.193
0.280
0.313
shot
Moirai
0.258
0.270
0.131
0.247
0.389
0.353
zero
Moira-MoE
0.260
0.279
0.196
0.316
0.684
0.496
shot
Moirai
0.294
0.296
0.144
0.268
0.554
0.413
Table 7: Comparison with pre-trained models on the covariate-informed forecasting
C.3 Full Results of Multivariate Forecasting
The full results of multivariate forecasting in Section 4.2 are listed in Table 8. We adopt the rolling forecasting approach (Liu et al. 2024c), where a model trained with predetermined input and forecasting lengths is evaluated with a longer forecasting length by iteratively integrating its output into the subsequent input. Following previous works (Liu et al. 2024c, 2025), we set the input length to 672, the training forecasting length to 96, the testing forecasting length to {96, 192, 336, 720}, and the patch length to 96.
Models
CITRAS
TimeXer
Timer-XL
iTrans.
Leddam
CARD
ModernTCN
FITS
DLinear
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTh1
96
0.355
0.395
0.391
0.425
0.364 0.397
0.387
0.418
0.371
0.406
0.373
0.403
0.372
0.400
0.378
0.401
0.369
0.400
192
0.386
0.416
0.420
0.443
0.405
0.424
0.416
0.437
0.397
0.422
0.402
0.422
0.405
0.421
0.410
0.420
0.405
0.422
336
0.403
0.429
0.441
0.459
0.427
0.439
0.434
0.450
0.417
0.435
0.424
0.436
0.428
0.437
0.431
0.433
0.435
0.445
720
0.412
0.445
0.460
0.483
0.439
0.459
0.447
0.473
0.441
0.463
0.459
0.472
0.446
0.468
0.441
0.451
0.493
0.508
Ave
0.389
0.421
0.428
0.453
0.409
0.430
0.421
0.445
0.406
0.431
0.414
0.433
0.413
0.432
0.415
0.426
0.426
0.444
ETTh2
96
0.293
0.358
0.281
0.350
0.301
0.354
0.307
0.365
0.275
0.345
0.282
0.345
0.271
0.340
0.277
0.342
0.286
0.354
192
0.351
0.395
0.339
0.391
0.360
0.396
0.376
0.406
0.330
0.381
0.338
0.381
0.327
0.378
0.334
0.377
0.344
0.393
336
0.372
0.416
0.371
0.419
0.382
0.419
0.416
0.435
0.355
0.402
0.361
0.401
0.351
0.399
0.358
0.398
0.369
0.416
720
0.419
0.453
0.422
0.459
0.443
0.468
0.432
0.456
0.394
0.434
0.380
0.423
0.396
0.433
0.387 0.426
0.408
0.454
Ave
0.359
0.405
0.353
0.405
0.372
0.409
0.383
0.415
0.339
0.391
0.340
0.388
0.336
0.388 0.339
0.386
0.352
0.404
ETTm1
96
0.282
0.340
0.300
0.358
0.297 0.347
0.312
0.367
0.303
0.351
0.305
0.351
0.311
0.360
0.308
0.351
0.307
0.350
192
0.329
0.369
0.342
0.381
0.344
0.376
0.351
0.390
0.339
0.372
0.336
0.370
0.347
0.380
0.338
0.367
0.337
0.368
336
0.365
0.391
0.383
0.402
0.382
0.399
0.389
0.412
0.370
0.392
0.365
0.387
0.381
0.400
0.367
0.384
0.366
0.386
720
0.423
0.425
0.450
0.436
0.450
0.437
0.458
0.449
0.423
0.422
0.415
0.416
0.438
0.433
0.418
0.413
0.418
0.418
Ave
0.350
0.382
0.368
0.394
0.368
0.390
0.377
0.404
0.358
0.384
0.355
0.381
0.369
0.393
0.358
0.379
0.357
0.380
ETTm2
96
0.174
0.260
0.176
0.263
0.180
0.263
0.183
0.272
0.163 0.253
0.162
0.252
0.178
0.269
0.164
0.255
0.166
0.259
192
0.231
0.299
0.235
0.301
0.244
0.306
0.241
0.310
0.218 0.292
0.214
0.287
0.235
0.307
0.219
0.293
0.222
0.300
336
0.280
0.333
0.287
0.335
0.299
0.344
0.296
0.346
0.271 0.328
0.263
0.319
0.288
0.340
0.272
0.328
0.278
0.339
720
0.365
0.389
0.368
0.386
0.377
0.397
0.393
0.409
0.363
0.385
0.351
0.375
0.366
0.388
0.360 0.382
0.379
0.405
Ave
0.262
0.320
0.266
0.321
0.275
0.327
0.278
0.334
0.254
0.315
0.247
0.308
0.267
0.326
0.254 0.314
0.262
0.326
ECL
96
0.128 0.222
0.131
0.231
0.127
0.219
0.133
0.229
0.130
0.224
0.129
0.225
0.161
0.270
0.142
0.244
0.138
0.238
192
0.149
0.242
0.149
0.248
0.145
0.236
0.158
0.258
0.148
0.241 0.147
0.242
0.172
0.281
0.156
0.256
0.152
0.251
336
0.164
0.259
0.168
0.268
0.159
0.252
0.168
0.262
0.164 0.257 0.164
0.259
0.183
0.292
0.173
0.272
0.167
0.268
720
0.199 0.288
0.214
0.309
0.187
0.277
0.205
0.294
0.202
0.291
0.202
0.292
0.216
0.321
0.213
0.304
0.203
0.302
Ave
0.160 0.253
0.166
0.264
0.155
0.246
0.164
0.258
0.161
0.253 0.160
0.255
0.183
0.291
0.171
0.269
0.165
0.265
Weather
96
0.152
0.205
0.150
0.203
0.157
0.205
0.174
0.225
0.149
0.201 0.147
0.204
0.153
0.210
0.145
0.197
0.169
0.229
192
0.200
0.249
0.192
0.242
0.206
0.250
0.227
0.268
0.194
0.243
0.191
0.244
0.199
0.250
0.186
0.237
0.211
0.268
336
0.253
0.291
0.240 0.280
0.259
0.291
0.290
0.309
0.242
0.283
0.238
0.282
0.251
0.290
0.241
0.277
0.258
0.306
720
0.325
0.342
0.309
0.329
0.337
0.344
0.374
0.360
0.314
0.335
0.309
0.334
0.323
0.341
0.342
0.347
0.320
0.362
Ave
0.233
0.272
0.223
0.263
0.240
0.273
0.266
0.291
0.225
0.265
0.221
0.266
0.232
0.273
0.229
0.265
0.239
0.291
Traffic
96
0.338
0.235
0.362
0.264
0.340 0.238
0.353
0.259
0.356
0.256
0.373
0.266
0.432
0.320
0.390
0.275
0.399
0.285
192
0.365
0.247
0.381
0.275
0.360
0.247
0.373
0.267
0.376
0.264
0.386
0.272
0.434
0.321
0.402
0.279
0.409
0.290
336
0.392
0.258
0.402
0.288
0.377
0.256
0.386
0.275
0.390
0.271
0.401
0.279
0.441
0.323
0.415
0.285
0.422
0.297
720
0.419
0.276
0.457
0.324
0.418
0.279
0.425
0.296
0.423
0.290
0.441
0.303
0.476
0.344
0.453
0.306
0.461
0.319
Ave
0.379
0.254
0.401
0.288
0.374
0.255
0.384
0.274
0.386
0.270
0.400
0.280
0.446
0.327
0.415
0.286
0.423
0.298
PEMS04
96
0.105
0.201
0.123
0.238
0.104
0.207
0.111
0.216
0.105
0.209
0.141
0.248
0.125
0.245
0.212
0.306
0.196
0.296
192
0.118
0.212
0.136
0.249
0.117 0.218
0.123
0.225
0.115
0.218
0.162
0.264
0.133
0.251
0.229
0.317
0.214
0.310
336
0.130
0.221
0.150
0.262
0.129 0.228
0.135
0.236
0.127
0.228
0.187
0.285
0.142
0.260
0.252
0.334
0.236
0.328
720
0.151
0.244
0.180
0.295
0.155 0.255
0.168
0.269
0.158
0.258
0.254
0.342
0.163
0.289
0.354
0.404
0.330
0.398
Ave
0.126
0.220
0.147
0.261
0.126
0.227
0.135
0.237
0.126
0.228
0.186
0.285
0.141
0.261
0.262
0.340
0.244
0.333
PEMS08
96
0.140
0.192
0.185
0.242
0.163 0.207
0.190
0.225
0.174
0.218
0.239
0.277
0.261
0.286
0.362
0.338
0.325
0.335
192
0.223
0.207
0.252
0.261
0.245 0.224
0.266
0.240
0.250
0.233
0.315
0.302
0.321
0.305
0.422
0.357
0.381
0.358
336
0.282
0.220
0.296
0.282
0.283 0.237
0.315
0.255
0.303
0.246
0.363
0.326
0.359
0.323
0.460
0.376
0.418
0.378
720
0.300
0.246
0.346
0.332
0.299
0.267
0.349
0.291
0.333
0.280
0.403
0.377
0.406
0.381
0.526
0.441
0.506
0.449
Ave
0.236
0.216
0.270
0.279
0.247 0.234
0.280
0.253
0.265
0.244
0.330
0.321
0.337
0.324
0.442
0.378
0.408
0.380
$1^{\text{st}}$
count
16
20
1
2
12
7
0
0
3
0
11
6
4
1
2
10
0
0
Table 8: Full results of the multivariate forecasting, where the input length is set to 672, and the forecasting length is set to {96, 192, 336, 720}. The evaluation protocol is the same as in Table 2. Results are cited from Timer-XL (Liu et al. 2025) if available; otherwise reproduced.
Appendix D Limitations and Future Work
In this paper, we explored effective ways to incorporate time-varying covariates into the forecasting process. We consider static covariates to be outside the scope of this paper and leave them for future work. Recently, numerous papers have tackled large-scale pre-training to achieve generalization performance in time series forecasting. Our model is based on a decoder-only Transformer trained through next-token prediction. This design is assumed to be suitable for large-scale pre-training, but its generalization capability has not yet been evaluated. In the future, we will explore the generalization potential of CITRAS in covariate-informed settings.